Two large, non-conducting plates with the indicated area charge densities are suspended 7.79 cm apart. You may treat each plate as an infinite sheet. How much electrostatic energy is stored in 4.54 cubic centimeter of the space in region A (left of both plates)?

What volume of the space in region B (between the plates) stores an equal amount of energy?

The plates are +81.3 uC/m^2 and +23.6 uC/m^2.
Answer #1

a)

Electric field in region A is

E_{A}=-rac{sigma _{1}}{2varepsilon _{o}}-rac{sigma _{2}}{2varepsilon _{o}}=-rac{1}{2 imes (8.8542 imes 10^{-12})}(81.3+23.6) imes 10^{-6}

EA=-5.9237*106 N/C

Electrostatic energy stored in region A is

U=rac{1}{2}varepsilon _{o}E^{2}V=rac{1}{2} imes (8.8542 imes 10^{-12})(-5.9237 imes 10^{6})^{2} imes (4.54 imes 10^{-6})

U=7.053*10-4 J

b)

Electric field in region B is

σ1 σ2 2, 2,-一 2 × (8.8542 × 10-12) (01.3-23.6) × 10 -6

EB=3.2583*106N/C

Since

U =-EÆgy ー× (8.8542 × 10-12)(3.2583 × 106)2 × VO

7.054 × 10-4-x(8.8542 × 10-12)(3.2583 × 106)2 × Vg

VB=1.5*10-5 m3=15 cm3